# Harm Amplification in Text-to-Image Models

### v2

@@ Abstract
Text-to-image (T2I) models have emerged as a significant advancement in generative AI; however, there exist safety concerns regarding their potential to produce harmful image outputs even when users input seemingly safe prompts. This phenomenon, where T2I models generate harmful representations that were not explicit in the input, poses a potentially greater risk than adversarial prompts, leaving users unintentionally exposed to harms. Our paper addresses this issue by formalizing a definition for this phenomenon which we term harm amplification. We further contribute to the field by developing a framework of methodologies to quantify harm amplification in which we consider the harm of the model output in the context of user input. We then empirically examine how to apply these different methodologies to simulate real-world deployment scenarios including a quantification of disparate impacts across genders resulting from harm amplification. Together, our work aims to offer researchers tools to comprehensively address safety challenges in T2I systems and contribute to the responsible deployment of generative AI models.
@@ Introduction
Generative text-to-image (T2I) systems allow users to create new image content in response to a text input, learning patterns and relationships from large-scale training data (Open AI 2023; Midjourney Inc. 2022; Stability AI 2022). As all data are reflections of the social world (Gitelman 2013), training data can mirror the harmful stereotypes and social inequalities that exist offline (Birhane et al. 2023; Birhane, Prabhu, and Kahembwe 2021; Kirk et al. 2021; Wang et al. 2022), which are subsequently replicated in generated imagery (Cho, Zala, and Bansal 2023). Responsible AI scholarship illustrates how content from T2I systems can reflect and amplify the harmful representations or stereotypes of groups that exist offline (Qu et al. 2023; Qadri et al. 2023; Rando et al. 2022b). Harms from generative AI — like all algorithmic systems — occur from the interplay of technical system components and social power dynamics (Shelby et al. 2023; Green and Viljoen 2020). Thus, effective harm reduction work requires practitioners to develop sociotechnical safety requirements, conduct evaluations, and implement strategies to proactively minimize harmful representations prior to launch (Rismani et al. 2023a; Dobbe 2022; Solaiman et al. 2023).
Responsible AI research has cultivated safety techniques to curb the generation of unsafe content including through reinforcement learning (Daniels-Koch and Freedman 2022), dataset remediation (Birhane et al. 2023), red teaming (Parrish et al. 2023; Rando et al. 2022b), model input and output filters (Hao et al. 2023), and techniques to suppress harmful content generation (Schramowski et al. 2023). These component safety techniques play a vital role in reducing the likelihood a model will generate harmful representations. However, these techniques do now eliminate all potential harms such as the exposure of unintentional contextualized harms in generative AI systems. Consider a scenario where a user requests an image of “people at a party” from a multimodal AI model, and images of women in bikinis dancing are outputted. The input prompt and output image in isolation may be considered safe, but when taken together in context of each other, they can expose unintentional harm (oversexualization of women) to the user.
This scenario underscores the complex challenges posed by multimodal systems. Harmful representations may result from computational, contextual, and compositional risks between text-image pairings with visual content introducing “additional shock factor[s]” (Weidinger et al. 2023, p. 13). While responsible AI scholarship acknowledges the amplification of harm as a key negative consequence of unsafe AI systems, including the “prospect of algorithmic systems exacerbating or scaling existing social inequalities” (Shelby et al. 2023, p. 733), a precise definition of multimodal safety harm amplification has yet to be introduced, to our knowledge.
To address these challenges, we formalize the concept of harm amplification in T2I systems and present a framework to quantify harm amplification. This framework employs three statistical methods that can be used in different contexts, including scenarios with: 1) large amounts of data and access to safety classifiers 2) limited data, but having access to safety classifiers, and 3) low resource settings with access only to open source models. In summary, we make the following research contributions:
@@
  • A framework that includes three methods for quantifying and measuring harm amplification instances in T2I models. These methods include using other pretrained models such as safety classifiers and CLIP (Radford et al. 2021) to determine whether harm amplification has occurred (i.e., when the output image is more harmful than the user input prompt).
  • An empirical examination of how to employ each methodology to simulate different real-world model deployment scenarios, which can dynamically inform practitioners’ safety frameworks. This includes a quantification of disparate impact across gendered outputs as a result of harm amplification.
In what follows, we first orient our work with respect to key concepts, including sociotechnical approaches to safety in T2I systems and stereotype amplification; followed by a description of our novel methodologies. We then detail our experiments and finish with a discussion of how empirically measuring harm amplification strengthens safety work on T2I systems, arguing for greater attention to harm amplification in responsible AI research and practice.
@@
This research lies at the intersection and builds on responsible AI research on harm, safety, and multimodal generative AI. The contributions of this research extend the work summarized, and highlight the need for a formalized definition of- and evaluating- harm amplification.
@@
Concerns about safety in AI systems encompass a range of diverse sociotechnical issues that directly or indirectly shape the well-being of people or the environment. As such, safety is contingent and contextual, and an emergent characteristic or property of AI systems (Leveson 2020). While approaches to safety in AI systems vary (Ahmed et al. 2023), a key concern of a systems approach is to identify system-level hazards (Dobbe 2022) and integrate safety considerations into the development process (Leveson 2016) to minimize the likelihood that harms will arise from system use or failure (Dobbe 2022). Within this approach, social and ethical safety considerations of T2I systems need to be identified by proactively examining the interaction between components of the AI system (e.g., technical capabilities, potential failures, context of use, intentions of users or other human actors) (Rismani et al. 2023a; Weidinger et al. 2023). In this way, safety considerations are reciprocally shaped by social ideas and dynamics that are encoded into AI systems (Dobbe 2022; Rismani et al. 2023b). This concept has been described by Ruha Benjamin (2019) as the “duplicity of technology,” in which AI systems often default to or replicate social power dynamics. Addressing this duplicity requires practitioners to understand social inequalities and how they may manifest in sociotechnical systems. Without intentional consideration and intervention, T2I systems can easily amplify harmful representations (Qadri et al. 2023; Bianchi et al. 2023).
AI Governance and Mitigating “Unsafe” Content
@@
An effective governance approach to cultivating “safe” AI systems is to develop clear safety requirements (Leveson 2020), with attention to the sociotechnical nature of AI systems (Dobbe 2022) and potential harms (Shelby et al. 2023). There is increasing recognition of representational harms in T2I systems (Cho, Zala, and Bansal 2023; Díaz et al. 2023; Bird, Ungless, and Kasirzadeh 2023), which are the socially constructed beliefs about different social groups that reinforce unjust hierarchies (Wang et al. 2022; Barocas et al. 2017; Barocas, Hardt, and Narayanan 2019), and manifest in different ways (Katzman et al. 2021). Representational harms are sociotechnical in nature (Shelby et al. 2023), as social beliefs about people, culture, and experiences are encoded into systems through training data and learned associations (Birhane, Prabhu, and Kahembwe 2021). As such, an increasingly common safety requirement (and social and regulatory expectation) (Solaiman 2023; Dias Oliva 2020) is reducing the likelihood that T2I systems will amplify harmful representations — such as demeaning gender, racial, disability, or cultural stereotypes (Qadri et al. 2023; Bianchi et al. 2023; Ungless, Ross, and Lauscher 2023; Naik and Nushi 2023; Bennett et al. 2021), particularly when model inputs are benign, unspecified, or do not contain qualifiers (i.e., “irrespective of gender”) (Bansal et al. 2022, p. 1359). It is necessary to address representational harms as they perpetuate normative narratives that adversely shape people’s sense of identity and belonging (e.g., (Bennett and Keyes 2020; Costanza-Chock 2018; Mengesha et al. 2021)) and algorithmic systems can scale the visibility of harmful representations reinforcing their social power (Phillips 2018).
Gender, Social Inequality, and Representation
@@
Gender inequalities are a persistent global problem, which can be exacerbated by AI technologies (Lütz 2023). Gender is a social construct through which a society defines expectations around masculinity and femininity, or outside this binary (Lorber 1996; West and Zimmerman 1987; West and Fenstermaker 1995). As a social construction, gender ideologies are created and changed to meet social, political, and economic needs (Martin 2004). They are an enduring major component structuring individual, interactional, and institutional power relations (Risman 2004). A dominant gender construct is the sexual objectification, where women are viewed as physical objects and valued for their sexual fulfilment (Fredrickson and Roberts 1997; Szymanski, Moffitt, and Carr 2011). Scholars have long highlighted how constructed ideas about gender difference are used to justify stratification and inequality (Lorber 1994), and that gender ideologies intersect with other axes of inequality, including race and ethnicity, socioeconomic status, sexuality, and disability among others. The concept of intersectionality was developed by feminist scholars of color to contest essentialized, white, middle-class, Western constructions of gender (e.g., (Anzaldúa 1987; Crenshaw 1991a, b; King 1988; Mohanty 2003)), revealing how gendered representations are informed by intersecting gender, race, and social class ideologies within a “matrix of domination” (Collins 1990). Computing scholars have examined gendered stereotypes, tropes, and patterns of over/under representation in algorithms (Buolamwini and Gebru 2018), datasets (Zhao et al. 2017; Meister et al. 2023; Garrido-Muñoz, Martínez-Santiago, and Montejo-Ráez 2023), and interfaces (Kay, Matuszek, and Munson 2015; Krieg et al. 2023), demonstrating persistent stereotype exaggeration (Kay, Matuszek, and Munson 2015). Gendered stereotypes affect the performance of real world applications, including online advertising (Sweeney 2013), toxic language detection (Park, Shin, and Fung 2018), machine translation and language technologies (Font and Costa-Jussà 2019; Dev et al. 2021; Vanmassenhove, Hardmeier, and Way 2018), search engines (Kopeinik et al. 2023; Kay, Matuszek, and Munson 2015; Albawardi and Jones 2023), and T2I systems (Bianchi et al. 2023; Naik and Nushi 2023).
@@
The representational harm phenomenon in T2I models has been clearly identified (Hao et al. 2023; Bird, Ungless, and Kasirzadeh 2023; Bianchi et al. 2023; Cho, Zala, and Bansal 2023; Ungless, Ross, and Lauscher 2023; Wolfe et al. 2023; Yu et al. 2022; Jain et al. 2021; Choi et al. 2020). Existing work often focuses on two areas: (1) identifying the kinds and types of harmful representations or stereotypes across different social attributes and axes of identity, and (2) evaluating and mitigating representational societal harms through measurement and technical intervention.
It is of note that stereotype amplification and harm amplification, although related, differ in key ways. Stereotype amplification in AI systems involves the reinforcement of stereotypes related to a group of people and does not necessarily pertain to safety. Conversely, harm amplification as we define it, occurs when AI systems exacerbates harms in the context of safety without necessarily targeting specific demographic groups. While stereotype amplification has been well studied, the formalization of harm amplification, to our knowledge, is novel.
Furthermore, our understanding of harm amplification is related, but distinct, from extant work on “bias amplification” examining how patterns of bias appear in training data, and are amplified statistically in model outputs (e.g., (Zhao et al. 2017; Wang and Russakovsky 2021; Zhao, Andrews, and Xiang 2023; Lloyd 2018; Taori and Hashimoto 2023; Mehrabi et al. 2021)). In this context, “bias amplification” reflects when a model makes “certain predictions at a higher rate for some groups than expected based on training-data statistics” (Hall et al. 2022a, p. 1), often focusing on binary classification models (e.g., (Leino et al. 2019; Wang et al. 2022; Hall et al. 2022b)) and correlations (Zhao et al. 2017). While bias amplification deals with statistical disparities in model outputs across groups, harm amplification encompasses a broader range of negative consequences not necessarily linked to a specific group.
Reinforcing Harmful Societal Representations
Early examination into social stereotype amplification focused on how Generative Adversarial Networks (GANs) exacerbate social biases along axes of gender and skin tone (Jain et al. 2021; Choi et al. 2020). Jain et al. (2021) demonstrate how generated images have a “distributional shift” away from real world data, in which generative AI systems exacerbate extant social inequalities and patterns of over/under-representation. The study of representational harms in T2I systems has extended beyond the relationships between training data and model outputs, to focus on the relationships between text prompts and resulting imagery. Bianchi and colleagues (2023) delineate and document many examples of the amplification of stereotypes in T2I systems, particularly with respect to racial and cultural stereotypes in unspecified text prompts. They show using Stable Diffusion, for example, that T2I systems amplify white racialized notions of beauty, in which “an attractive person” generates faces similar to a “white ideal” (p. 4); in another example, they illustrate that a prompt for “a terrorist” generates brown faces with dark hair and beards, reflecting demeaning stereotypes at the intersections of race and religion (p. 3). Moving beyond representations of people, the authors also show that generating images of everyday objects (e.g., “a house”) reflects Western and middle-class stereotypes (p. 3). Adding specific targeted identity terms (i.e., writing a specified prompt) did not completely mitigate the stereotypes and bias in T2I imagery. Qadri et al. (2023, p. 506) examine the cultural limitations of T2I systems by soliciting South Asian community experts to generate text prompts and annotate the resulting imagery, revealing T2I models “can participate in and scale such existing harmful media regimes of representation and amplify experiences of socio-cultural marginalization.” Similarly, Naik et al. (2023) explore exacerbation of social inequalities along axes of (perceived) gender, age, race, and geographical location, showing state-of-the-art T2I systems amplify real world inequities when compared to U.S. Bureau of Labor Statistics. These studies find that explicitly adding identity terms to unspecified prompts (e.g., “female security guard” (Naik and Nushi 2023)) does not effectively mitigate the presence of stereotypes in generated imagery (Bianchi et al. 2023).
@@
Evaluating Representational Societal Biases
A parallel body of research focuses not just on identifying the presence and kinds of harmful stereotypes that T2I systems can generate, but also evaluating them. Luccioni and colleagues (2023) examine whether representational societal harms in each modality amplify or compound one another by evaluating images generated from two synthetic text prompt datasets: one with gender and ethnicity terms and one with profession identifiers. The authors analyze representations in T2I imagery, assessing the correlations between social attributes in text markers and the image space, and comparing these to U.S. labor demographics, finding consistent under-representation of marginalized identities. Wang et al. (2023) employ the Text-to-Image Association Test to measure and quantify the presence of societal bias in T2I models while measuring stereotype amplification in occupations across genders. Another line of work has been to detect and measure social biases in downstream applications. Perera et al. (2023) explore social bias with respect to gender, race, and age in the diffusion-based face generation models, finding that diffusion models exacerbate patterns of over- and under-representation, while Saravanan and colleagues (2023) conducts an intersectional analysis of racialized gender stereotypes and professions to support evaluation of T2I models. While a growing body of research is exploring methods to measure and evaluate social stereotypes, safety harm amplification has yet to be formally defined. Furthermore, a comprehensive examination of harm representation across social groups will further support sociotechnical safety work and contribute to the social and ethical evaluation of T2I systems.
@@
Our paper focuses on examining unsafe amplifications of harms in a specific set of generative AI: T2I systems. First, we define “unsafe” content in T2I systems as the presence of demeaning, contextually inappropriate, or offensive content in generated images as outlined in safety requirements. What could be considered “unsafe” or harmful is highly contextual, and needs to be defined within the broader sociotechnical dynamics in which the system is deployed, including system capabilities, context of use, user audiences, and social or regulatory expectations (Rismani et al. 2023a; Dobbe 2022; Leveson 2020). In practice, what constitutes “unsafe” or harmful content may be informed by social, regulatory, and use case expectations; and should be clearly defined with respect to the AI system as a whole. For illustrative purposes in this paper, we explore two specific kinds of harmful representations to illustrate how harm amplification can be quantified in T2I systems, following the definitions of Hao et al. (2023):
@@
After harms and safety requirements have been established for a specific T2I system and context, harm amplification can be defined as occurring when the output generated image is more harmful than the input text prompt. In this sense, amplification reflects when a T2I system increases the severity of the defined harms without being explicitly prompted to do so (e.g., sexually objectifying women in generated imagery). Thus, the notion of harm amplification we propose is concerned with understanding the relationships between the text input and resulting imagery. In this paper, we choose not to articulate a specific threshold for what constitutes as harm amplification as we believe this should be decided within a specific product or technology context with input from a wider range of stakeholders. Instead, we offer a methodological framework that demonstrates how to derive thresholds in various deployment scenarios and show in our experimental section how to employ this framework on an example dataset.
@@
Quantifying harm amplification is a crucial component in assessing “safety” in T2I systems with respect to defined unsafe content. In this section, we introduce three statistical methods designed to measure harm amplification. Leveraging safety classifiers and embedding models, our approach aims to understand how the harms of the generated output images are amplified in the context of the user input prompts. Each of our methods has strengths and limitations that makes them well-suited for specific cases and available resources. This tailored approach ensures that our methods can be used across diverse contexts contributing to a more robust assessment of harm amplification.

### v3

@@ Abstract
Text-to-image (T2I) models have emerged as a significant advancement in generative AI; however, there exist safety concerns regarding their potential to produce harmful image outputs even when users input seemingly safe prompts. This phenomenon, where T2I models generate harmful representations that were not explicit in the input prompt, poses a potentially greater risk than adversarial prompts, leaving users unintentionally exposed to harms. Our paper addresses this issue by formalizing a definition for this phenomenon which we term harm amplification. We further contribute to the field by developing a framework of methodologies to quantify harm amplification in which we consider the harm of the model output in the context of user input. We then empirically examine how to apply these different methodologies to simulate real-world deployment scenarios including a quantification of disparate impacts across genders resulting from harm amplification. Additionally, we employ explainability techniques to understand mechanisms driving harm amplification and how specific input elements contribute to harmful outputs. Together, our work aims to offer researchers tools to comprehensively address safety challenges in T2I systems and contribute to the responsible deployment of generative AI models.
@@ Introduction
Generative text-to-image (T2I) systems allow users to create new image content in response to a text input prompt. These systems learn patterns and relationships from their large-scale training data (Open AI 2023; Midjourney Inc. 2022; Stability AI 2022). However, they frequently reflect harmful stereotypes and social inequalities that are embedded in their training data (Birhane et al. 2023; Birhane, Prabhu, and Kahembwe 2021; Kirk et al. 2021; Wang et al. 2022), which are subsequently reproduced in generated imagery (Cho, Zala, and Bansal 2023). The interaction between these technical systems and existing social power dynamics in the world (Shelby et al. 2023; Green and Viljoen 2020) can further perpetuate and amplify the harmful representations of marginalized groups (Qu et al. 2023; Qadri et al. 2023; Rando et al. 2022b).
Responsible AI research has often approached the study of harmful content in generative AI systems using adversarial prompts and red teaming (Parrish et al. 2023; Rando et al. 2022b). While these methods are valuable, they overlook a critical aspect of safety - unintentional contextual harms arising from seemingly benign inputs. For example, a user requesting an image of a “black gay man” may inadvertantly receive sexualized images from a generative AI system, exposing them to unintentional harm through oversexualization which may further perpetuate harmful societal biases. Even with safety filters in place (Hao et al. 2023), users may experience frustration and distress when safe prompts are blocked without explanation.
This scenario underscores the complex challenges posed by multimodal systems where harmful representations can result from computational, contextual, and compositional risks. While responsible AI scholarship acknowledges the amplification of harm as a key negative consequence of unsafe AI systems, including the “prospect of algorithmic systems exacerbating or scaling existing social inequalities” (Shelby et al. 2023, p. 733), there is currently no precise definition of multimodal safety harm amplification.
To address these challenges, we provide the first formalization of harm amplification and make the following research contributions:
@@
  • A framework that includes three methods for quantifying and measuring harm amplification instances in T2I models.
  • An empirical examination of patterns in harm amplification including quantifying its disparate impact across genders and applying explainability techniques to identify aspects of the user input that contribute to harmful output generation.
In what follows, we first orient our work with respect to key concepts, including sociotechnical approaches to safety in T2I systems and stereotype amplification; followed by a description of our novel methodologies and their efficacy in measuring harm amplification. We then examine patterns of harm amplification within our data and finish with a discussion of how empirically measuring harm amplification strengthens safety work on T2I systems, arguing for greater attention to harm amplification in responsible AI research and practice.
@@
This research builds on existing responsible AI literature addressing harm, safety, and multimodal generative AI, specifically T2I systems. Our contributions extend the work summarized, and highlight the need for a formalized definition of harm amplification and methods for evaluating it.
@@
AI safety encompass a range of diverse sociotechnical issues that impact the well-being of people and the environment. A systems approach to AI safety identifies hazards (Dobbe 2022) and integrates safety considerations into the development process (Leveson 2016) to minimize potential harms (Dobbe 2022). Addressing social and ethical safety in T2I systems requires examining the interactions between AI components and social dynamics (Rismani et al. 2023; Weidinger et al. 2023). AI systems can replicate social dynamics (Benjamin 2019) or even amplify harmful representations that perpetuate societal biases (Qadri et al. 2023; Bianchi et al. 2023).
@@
An effective governance approach to cultivating “safe” AI systems is to develop clear safety requirements (Leveson 2020), with attention to the sociotechnical nature of AI systems (Dobbe 2022) and potential harms (Shelby et al. 2023). There is increasing recognition of representational harms in T2I systems (Cho, Zala, and Bansal 2023; Díaz et al. 2023; Bird, Ungless, and Kasirzadeh 2023), which are the socially constructed beliefs about different social groups that reinforce unjust hierarchies (Wang et al. 2022; Barocas et al. 2017; Barocas, Hardt, and Narayanan 2019), and manifest in different ways (Katzman et al. 2021). Representational harms are sociotechnical in nature (Shelby et al. 2023), as social beliefs about people, culture, and experiences are encoded into systems through training data and learned associations (Birhane, Prabhu, and Kahembwe 2021).
@@
Gender inequalities, exacerbated by AI technologies, illustrate a significant global challenge (Lütz 2023) and exemplify a specific instance of representational harms. Gender, a social construct, shapes expectations around masculinity, and femininity, and beyond (Lorber 1996; West and Zimmerman 1987; West and Fenstermaker 1995). Gender ideologies adapted to meet various social needs (Martin 2004) and are a major component in structuring individual, interactional, and institutional power relations (Risman 2004). A dominant gender construct is sexual objectification, where women are viewed as physical objects and valued for their sexual fulfilment (Fredrickson and Roberts 1997; Szymanski, Moffitt, and Carr 2011). This along with other gender stereotypes affect the performance of real world applications, including online advertising (Sweeney 2013), toxic language detection (Park, Shin, and Fung 2018), machine translation and language technologies (Font and Costa-Jussà 2019; Dev et al. 2021; Vanmassenhove, Hardmeier, and Way 2018), search engines (Kopeinik et al. 2023; Kay, Matuszek, and Munson 2015; Albawardi and Jones 2023), and T2I systems (Bianchi et al. 2023; Naik and Nushi 2023).
@@
Early examination into social stereotype amplification focused on how Generative Adversarial Networks (GANs) exacerbate social biases along axes of gender and skin tone (Jain et al. 2021; Choi et al. 2020) demonstrating ways in which generative AI systems amplify extant social inequalities and patterns of over/under-representation. Bianchi and colleagues (2023) delineate and document many examples of the amplification of stereotypes in T2I systems, particularly with respect to racial and cultural stereotypes in unspecified text prompts. They show using Stable Diffusion, that T2I systems amplify white racialized notions of beauty, in which “an attractive person” generates faces similar to a “white ideal” (p. 4) and “a terrorist” generates brown faces with dark hair and beards. Qadri et al. (2023, p. 506) examine the cultural limitations of T2I systems by soliciting South Asian community experts to generate text prompts and annotate the resulting imagery, revealing T2I models “can participate in and scale such existing harmful media regimes of representation and amplify experiences of socio-cultural marginalization.” Similarly, Naik et al. (2023) explore exacerbation of social inequalities along axes of (perceived) gender, age, race, and geographical location, showing state-of-the-art T2I systems amplify real world inequities when compared to U.S. Bureau of Labor Statistics.
@@
It is of note that stereotype amplification and harm amplification, although related, differ in key ways. Stereotype amplification in AI systems involves the reinforcement of stereotypes related to a group of people and does not necessarily pertain to safety. Conversely, harm amplification as we define it, occurs when AI systems exacerbates harms in the context of safety without necessarily targeting specific demographic groups. While stereotype amplification has been well studied, the formalization of harm amplification is novel. Furthermore, our understanding of harm amplification is related, but distinct, from extant work on bias amplification examining how patterns of bias appear in training data, and are amplified statistically in model outputs (Zhao et al. 2017; Wang and Russakovsky 2021; Zhao, Andrews, and Xiang 2023; Lloyd 2018; Taori and Hashimoto 2023; Mehrabi et al. 2021). In this context, bias amplification reflects when a model makes “certain predictions at a higher rate for some groups than expected based on training-data statistics” (Hall et al. 2022a, p. 1), often focusing on binary classification models (e.g., (Leino et al. 2019; Wang et al. 2022; Hall et al. 2022b)). While bias amplification typically addresses demographic disparities in classifier accuracy, harm amplification extends beyond predication accuracy and statistical disparities, aiming to understand how the generated model output’s safety relates to the user input.
@@
Our paper focuses on examining unsafe amplifications of harms in a specific set of generative AI - T2I systems. First, we define “unsafe” content in T2I systems as the presence of demeaning, contextually inappropriate, or offensive content in generated images as outlined in safety requirements. What could be considered “unsafe” or harmful is highly contextual, and needs to be defined within the broader sociotechnical dynamics in which the system is deployed, including system capabilities, context of use, user audiences, and social or regulatory expectations (Rismani et al. 2023; Dobbe 2022; Leveson 2020). For illustrative purposes in this paper, we explore two specific kinds of harmful representations to illustrate how harm amplification can be quantified in T2I systems, following the definitions of Hao et al. (2023):
@@
After harms and safety requirements have been established for a specific T2I system and context, we propose the following definition for harm amplification. Let $T$ be the space of all possible input prompts, $I$ be the space of all generated images, and $M$ be a text-to-image model that maps input prompts to generated images. Let $H$ be a harm function that measures the level of harm of the input or output of the model. For a given input prompt $t\in T$ and its corresponding generated image $i=M(t)$, we define harm amplification as occurring when:
$H(i)>H(t)+\tau$ (1)
where
  • $H(i)$ is the harm level of the output image alone
  • $H(t)$ is the harm level of the input prompt alone
  • $\tau$ represents a predefined threshold parameter
Thus, harm amplification can be defined as occurring when the output generated image is more harmful than the input text prompt. In this sense, amplification reflects when a T2I system increases the severity of the defined harms without being explicitly prompted to do so (e.g., sexually objectifying women in generated imagery). Thus, the notion of harm amplification we propose is concerned with understanding the relationships between the text input and resulting imagery. In this paper, we choose not to articulate a specific threshold for what constitutes as harm amplification as we believe this should be decided within a specific product or technology context with input from a wider range of stakeholders. Instead, we offer a methodological framework that demonstrates how to derive thresholds in various deployment scenarios and show how to employ this framework on an example dataset.
@@
In this section, we introduce three methods to quantify harm amplification. We then evaluate the efficacy of these methods on the human annotated Nibbler evaluation dataset (Quaye et al. 2024).
Methods
Quantifying harm amplification is a crucial component in assessing safety in T2I systems with respect to defined unsafe content. In this section, we introduce three statistical methods designed to measure harm amplification. Leveraging safety classifiers and embedding models, our approach aims to understand how the harms of the generated output images are amplified in the context of the user input prompts. Each of our methods has strengths and limitations that makes them well-suited for specific cases and available resources. This tailored approach ensures that our methods can be used across diverse contexts contributing to a more robust assessment of harm amplification.